# Tensor renormalization group approach to critical phenomena via symmetry-twisted partition functions

### v1

@@ IV Numerical Results
Figure 1: $Z_{-1}/Z_{1}$ in the 2D Ising model. The vertical dashed line indicates the exact critical point $T_{c}=2/(\log(1+\sqrt{2}))$. The horizontal dashed line denotes the exact value of $Z_{-1}/Z_{1}$ for the 2D Ising CFT, which is given by Eq. (IV.2). The computation is done by the BTRG with the bond dimension $D_{\rm BTRG}=64$.
The partition function of the Ising model can be expressed as Eq. (III.2) with the initial bond dimension 2, and we compute its tensor contraction using the BTRG with the bond dimension cutoff $D_{\rm BTRG}=64$. Figure 1 shows the ratio of the twisted partition functions $Z_{-1}(L\times L)/Z_{1}(L\times L)$ as a function of temperature for various volumes. As the system size $L$ increases, the behavior of $Z_{-1}/Z_{1}$ exhibits a pronounced distinction across the critical point $T_{c}$: it approaches unity in the symmetric phase, while tending toward zero in the symmetry-broken phase, as expected.
@@
$\left.\frac{Z_{-1}(L_{1}\times L_{2})}{Z_{1}(L_{1}\times L_{2})}\right|_{T=T_{c}}=\frac{|\theta_{3}(0,\tau)|+|\theta_{4}(0,\tau)|-|\theta_{2}(0,\tau)|}{|\theta_{3}(0,\tau)|+|\theta_{4}(0,\tau)|+|\theta_{2}(0,\tau)|},$ (IV.2)
where $\theta_{2}$, $\theta_{3}$, and $\theta_{4}$ denote Jacobi’s theta functions, and $\tau={\rm i}L_{2}/L_{1}$ when we consider the model on the $L_{1}\times L_{2}$ square lattice. In Fig. 1, we set $\tau=\mathrm{i}$ as $L_{1}=L_{2}=L$ for the horizontal red dashed line. The scale-invariant point in the numerical results shows the nice consistency with the theoretical value of $Z_{-1}/Z_{1}$ in Eq. (IV.2), which validates the efficacy of our approach.
@@
$S=-\beta\sum_{x}\sum_{\mu=1}^{d}\cos(\theta_{x+\hat{\mu}}-\theta_{x}),$ (IV.3)
@@
$(T_{x})_{m_{1}\ldots m_{d},m^{\prime}_{1}\ldots m^{\prime}_{d}}=\sqrt{I_{m_{1}}(\beta)\cdots I_{m_{d}}(\beta)I_{m^{\prime}_{1}}(\beta)\cdots I_{m^{\prime}_{d}}(\beta)}\,\delta_{m_{1}+\cdots+m_{d},m^{\prime}_{1}+\cdots+m^{\prime}_{d}}$ (IV.4)
@@
at each lattice site $n$. The finite bond dimensional description of Eq. (IV.4) is obtained by truncating the character expansion, $\mathrm{e}^{\beta\cos\Theta}\simeq\sum_{m=-K}^{K}I_{m}(\beta)\mathrm{e}^{\mathrm{i}m\Theta}$, which results in the tensor network representation of the partition function with bond dimension $2K+1$ [48]. In the range of $\beta$ considered in this study, taking $K\geq 40$ already yields $I_{K}(\beta)/I_{0}(\beta)\lesssim 10^{-23}$, and the finite-$K$ effect is negligible. We note that the finite bond-dimension cutoff necessarily suffers from the mild violation of the $U(1)$ symmetry, as it gives the upper and lower bounds for the $U(1)$ charge sectors, which violates the $U(1)$ conservation law: While the finite truncation preserves the global $U(1)$ invariance of the partition function, it violates the Ward-Takahashi identity when including the local operators. When we consider the twisted partition function, we take the twist angles quantized in $2\pi/N$ for some integer $N$, and thus we require that the truncations and the tensor decompositions are consistent with the charge conservation in mod $N$ by introducing the $N$ block structure.
@@
Let us consider the 3D $O(2)$ model in this subsection. For the twisted partition functions, we take the $\mathbb{Z}_{2}$ subgroup of the $U(1)$ symmetry, and the twist angle is set as $\alpha=\pi$. For the 3D $O(2)$ model, the continuous symmetry is expected to be spontaneously broken completely as $U(1)\to\{1\}$ at low temperatures, and the second-order phase transition takes place at some critical temperature $T_{c}$. In the low temperature limit, $\beta=1/T\gg 1$, we can take the quadratic approximation of the lattice action (IV.3) to obtain the continuum effective Lagrangian (II.12) of the $U(1)$ SSB phase with $v^{2}=\beta/a^{d-2}=\beta/a$ with $v$ as defined in  (II.12), where $a$ is the lattice spacing. This suggests the $\mathbb{Z}_{2}$ twist does a sufficient job as the order parameter. The 3D $O(2)$ universality class has been extensively investigated in Monte Carlo simulations (see, for example, Refs. [32, 60, 61]). For the conformal bootstrap analysis about the 3D $O(2)$ criticality, see Refs. [62, 63]. The application of the TRG approach to the $O(2)$ model at finite density has also been reported [64].
@@
Figure 4: $Z_{\alpha=\pi}/Z_{0}$ in the 3D $O(2)$ model. The vertical dashed line indicates the critical point $T_{c}^{\mathrm{(MC)}}=2.2018441(5)$ obtained by the Monte Carlo simulation in Ref. [61]. The gray band denotes the value of $Z_{\alpha=\pi}/Z_{\alpha=0}$ at the critical point from the Monte Carlo simulation with $L=2^{4}$ [65]. The twisted partition function gives a clear signal for the $U(1)$ symmetry breaking. The computation is done by the ATRG with the bond dimension $D_{\rm ATRG}=96$.
@@
We can also see an emergent scale-invariant point at the phase transition $T=T_{c}$, which suggests that the phase transition is of the second order. The vertical red dotted line of Fig. 4 at $T_{c}^{(\mathrm{MC})}=2.2018441(5)$ is the critical temperature determined by the Monte Carlo simulation in Ref. [61], which is consistent with our numerical result with the ATRG. In addition, the gray band in Fig. 4 denotes the resulting $Z_{\alpha=\pi}/Z_{\alpha=0}$ at the critical point from another Monte Carlo simulation with $L=2^{4}$ [65], which again looks consistent with our result.
@@
Let us move on to the 2D classical $O(2)$ model, (IV.3). In contrast to dimensions three and higher, the Mermin–Wagner–Coleman theorem [36, 37] prohibits the presence of the massless Nambu–Goldstone particle with the linear dispersion in two dimensions. Thus, the $U(1)$ symmetry does not get broken even at low temperatures, while the system is in the superfluid phase, showing the algebraic long-range order. The effective Lagrangian of this superfluid phase is given by (II.12) when the vortex excitations are irrelevant.
@@
$\Upsilon_{\alpha}(L_{1}\times L_{2})=-T\frac{2}{\alpha^{2}}\frac{L_{2}}{L_{1}}\log\frac{Z_{\alpha}(L_{1}\times L_{2})}{Z_{0}(L_{1}\times L_{2})},$ (IV.10)
@@
$\Upsilon_{\alpha}(L_{1}\times L_{2})=\frac{T}{2\pi R_{\mathrm{ren}}^{2}}+O(\mathrm{e}^{-\#\frac{L_{1}}{L_{2}}})$ (IV.11)
@@
$\Upsilon_{\pi}(L_{1}\times L_{2})=\frac{T}{2\pi R_{\mathrm{ren}}^{2}}-\frac{2T}{\pi^{2}}\frac{L_{2}}{L_{1}}\ln 2+O(\mathrm{e}^{-\#\frac{L_{1}}{L_{2}}}).$ (IV.12)
@@
Equation (IV.11) shows that the helicity modulus $\Upsilon_{\alpha}$ dictates the renormalized compact boson radius directly for $0<\alpha\ll\pi$ when we can neglect the higher-winding contributions of $O(\mathrm{e}^{-\#L_{1}/L_{2}})$. This approximation becomes valid when $L_{1}/L_{2}\gg 1$. [^7] Ideally, it would be desired to take the limit $L_{1}/L_{2}\to\infty$ for completely eliminating higher-winding contributions. However, we should note that we compute $\Upsilon_{\alpha}$ out of $Z_{\alpha}/Z_{0}$, and this partition-function ratio itself becomes exponentially small as $L_{1}/L_{2}\to\infty$ according to (II.15). Thus, we should vary the ratio $L_{1}/L_{2}$ within the regime where $Z_{\alpha}/Z_{0}$ is computable with good accuracy. Moreover, as the higher-winding contributions are exponentially suppressed, we do not need to take extremely large $L_{1}/L_{2}$ in practice. This relationship becomes important when we apply the Nelson–Kosterlitz criterion [70] to obtain the BKT transition temperature $T_{\mathrm{BKT}}$, because it compares the helicity modulus $\Upsilon_{\alpha}$ with $2T/\pi$ to identify the marginality crossing point $R_{\mathrm{ren}}^{2}=1/4$. We note that this aspect-ratio dependence exists also for the helicity modulus defined by the local correlation function (or $\Upsilon_{\alpha}$ with the $\alpha\to 0$ limit), and see Refs. [71, 72], which study the aspect-ratio dependence of the helicity modulus. Although these higher-winding contributions turn out to be negligibly small within the numerical accuracy of this paper, [^8] For the $L\times L$ lattice, its contribution is already $O(10^{-4})$, and it becomes $O(10^{-9})$ for the $2L\times L$ lattice. we perform the computation not only at the $L\times L$ lattice but also at the $2L\times L$ lattice to suppress them.
@@
Figure 7 shows the helicity modulus $\Upsilon_{\alpha}(L\times L)$ evaluated at several twist angles, $\alpha=\pi,\pi/2,\pi/4,\pi/6$, for fixed $L=32$. We observe that the twist-angle dependence is quite tiny for $\Upsilon_{\alpha=\pi/2,\pi/4,\pi/6}$, which is consistent with the theoretical formula (IV.11) as its $\alpha$-dependence appears only in the subleading higher-winding contributions $O(\mathrm{e}^{-\#L_{1}/L_{2}})$. Furthermore, $\Upsilon_{\alpha=\pi}$ is significantly lower than the others on the low-temperature regions, which also matches with the theoretical expectation (IV.12). These helicity moduli at finite $\alpha$, obtained directly from the twisted partition functions, are found to be consistent with the recent Monte Carlo simulation results in Ref. [73].
@@
Figure 9: $T^{*}(L)$ as a function of $1/(\ln bL)^{2}$ for varying system sizes with the $2L\times L$ lattices: (a) $\alpha=\pi$, (b) $\alpha=\pi/6$. For $\alpha=\pi$, we determined $T^{*}(L)$ by using the modified Nelson–Kosterlitz criterion based on (IV.12). The results of different volumes follow the universal curve as a function of $1/(\ln bL)^{2}$, and we extract the BKT transition temperature by extrapolating $L\to\infty$.
Figure 9 shows $T^{*}(L)$ against $1/(\ln bL)^{2}$ at $\alpha=\pi$ and $\alpha=\pi/6$ evaluated on the $2L\times L$ lattices. When $\alpha=\pi$, $T^{*}(L)$ is determined based on Eq. (IV.12), not on the standard Nelson–Kosterlitz formula (IV.11), while the standard one is used for $\alpha=\pi/6$. In both cases, we simply neglect the correction from the higher-winding contributions as they are sufficiently small for the $2L\times L$ lattice with the current precision. The result is
$T_{\mathrm{BKT}}=\left\{\begin{array}[]{cl}0.8928(2)&(\text{from the $\alpha=\pi$ TRG}),\\ 0.8927(5)&(\text{from the $\alpha=\pi/6$ TRG}).\end{array}\right$. (IV.16)
@@ V Conclusions and Discussions
One of the potentially interesting applications is to consider the generalized $O(2)$ model, which extends the lattice action (IV.3) of the usual $O(2)$ model as [81, 82]

### v2

@@ IV Numerical Results
Figure 1: $Z_{-1}/Z_{1}$ in the 2D Ising model. The vertical dashed line indicates the exact critical point $T_{c}=2/(\log(1+\sqrt{2}))$. The horizontal dashed line denotes the exact value of $Z_{-1}/Z_{1}$ for the 2D Ising CFT, which is given by Eq. (IV.3). The computation is done by the BTRG with the bond dimension $D_{\rm BTRG}=64$.
The partition function of the Ising model can be expressed as Eq. (III.2) with the initial bond dimension 2, as
$(T_{x})_{m_{1}m_{2},m^{\prime}_{1}m^{\prime}_{2}}=\sqrt{I_{m_{1}}(\beta)I_{m_{2}}(\beta)I_{m^{\prime}_{1}}(\beta)I_{m^{\prime}_{2}}(\beta)}\,\delta_{m_{1}+m_{2},m^{\prime}_{1}+m^{\prime}_{2}},$ (IV.2)
where $I_{0}(\beta)=\cosh\beta$ and $I_{1}(\beta)=\sinh\beta$, and we compute its tensor contraction using the BTRG with the bond dimension cutoff $D_{\rm BTRG}=64$. Figure 1 shows the ratio of the twisted partition functions $Z_{-1}(L\times L)/Z_{1}(L\times L)$ as a function of temperature for various volumes. As the system size $L$ increases, the behavior of $Z_{-1}/Z_{1}$ exhibits a pronounced distinction across the critical point $T_{c}$: it approaches unity in the symmetric phase, while tending toward zero in the symmetry-broken phase, as expected.
@@
$\left.\frac{Z_{-1}(L_{1}\times L_{2})}{Z_{1}(L_{1}\times L_{2})}\right|_{T=T_{c}}=\frac{|\theta_{3}(0,\tau)|+|\theta_{4}(0,\tau)|-|\theta_{2}(0,\tau)|}{|\theta_{3}(0,\tau)|+|\theta_{4}(0,\tau)|+|\theta_{2}(0,\tau)|},$ (IV.3)
where $\theta_{2}$, $\theta_{3}$, and $\theta_{4}$ denote Jacobi’s theta functions, and $\tau={\rm i}L_{2}/L_{1}$ when we consider the model on the $L_{1}\times L_{2}$ square lattice. In Fig. 1, we set $\tau=\mathrm{i}$ as $L_{1}=L_{2}=L$ for the horizontal red dashed line. The scale-invariant point in the numerical results shows the nice consistency with the theoretical value of $Z_{-1}/Z_{1}$ in Eq. (IV.3), which validates the efficacy of our approach.
@@
$S=-\beta\sum_{x}\sum_{\mu=1}^{d}\cos(\theta_{x+\hat{\mu}}-\theta_{x}),$ (IV.4)
@@
$(T_{x})_{m_{1}\ldots m_{d},m^{\prime}_{1}\ldots m^{\prime}_{d}}=\sqrt{I_{m_{1}}(\beta)\cdots I_{m_{d}}(\beta)I_{m^{\prime}_{1}}(\beta)\cdots I_{m^{\prime}_{d}}(\beta)}\,\delta_{m_{1}+\cdots+m_{d},m^{\prime}_{1}+\cdots+m^{\prime}_{d}}$ (IV.5)
@@
at each lattice site $n$. The finite bond dimensional description of Eq. (IV.5) is obtained by truncating the character expansion, $\mathrm{e}^{\beta\cos\Theta}\simeq\sum_{m=-K}^{K}I_{m}(\beta)\mathrm{e}^{\mathrm{i}m\Theta}$, which results in the tensor network representation of the partition function with bond dimension $2K+1$ [48]. In the range of $\beta$ considered in this study, taking $K\geq 40$ already yields $I_{K}(\beta)/I_{0}(\beta)\lesssim 10^{-23}$, and the finite-$K$ effect is negligible. We note that the finite bond-dimension cutoff necessarily suffers from the mild violation of the $U(1)$ symmetry, as it gives the upper and lower bounds for the $U(1)$ charge sectors, which violates the $U(1)$ conservation law: While the finite truncation preserves the global $U(1)$ invariance of the partition function, it violates the Ward-Takahashi identity when including the local operators. When we consider the twisted partition function, we take the twist angles quantized in $2\pi/N$ for some integer $N$, and thus we require that the truncations and the tensor decompositions are consistent with the charge conservation in mod $N$ by introducing the $N$ block structure.
@@
Let us consider the 3D $O(2)$ model in this subsection. For the twisted partition functions, we take the $\mathbb{Z}_{2}$ subgroup of the $U(1)$ symmetry, and the twist angle is set as $\alpha=\pi$. For the 3D $O(2)$ model, the continuous symmetry is expected to be spontaneously broken completely as $U(1)\to\{1\}$ at low temperatures, and the second-order phase transition takes place at some critical temperature $T_{c}$. In the low temperature limit, $\beta=1/T\gg 1$, we can take the quadratic approximation of the lattice action (IV.4) to obtain the continuum effective Lagrangian (II.12) of the $U(1)$ SSB phase with $v^{2}=\beta/a^{d-2}=\beta/a$ with $v$ as defined in  (II.12), where $a$ is the lattice spacing. This suggests the $\mathbb{Z}_{2}$ twist does a sufficient job as the order parameter. The 3D $O(2)$ universality class has been extensively investigated in Monte Carlo simulations (see, for example, Refs. [32, 60, 61]). For the conformal bootstrap analysis about the 3D $O(2)$ criticality, see Refs. [62, 63]. The application of the TRG approach to the $O(2)$ model at finite density has also been reported [64].
@@
Figure 4: $Z_{\alpha=\pi}/Z_{0}$ in the 3D $O(2)$ model. The vertical dashed line indicates the critical point $T_{c}^{\mathrm{(MC)}}=2.2018441(5)$ obtained by the Monte Carlo simulation in Ref. [61]. The brown band denotes the value of $Z_{\alpha=\pi}/Z_{\alpha=0}$ at the critical point from the Monte Carlo simulation with $L=2^{4}$ [65]. The twisted partition function gives a clear signal for the $U(1)$ symmetry breaking. The computation is done by the ATRG with the bond dimension $D_{\rm ATRG}=96$.
@@
We can also see an emergent scale-invariant point at the phase transition $T=T_{c}$, which suggests that the phase transition is of the second order. The vertical red dotted line of Fig. 4 at $T_{c}^{(\mathrm{MC})}=2.2018441(5)$ is the critical temperature determined by the Monte Carlo simulation in Ref. [61], which is consistent with our numerical result with the ATRG. In addition, the brown band in Fig. 4 denotes the resulting $Z_{\alpha=\pi}/Z_{\alpha=0}$ at the critical point from another Monte Carlo simulation with $L=2^{4}$ [65], which again looks consistent with our result.
@@
Let us move on to the 2D classical $O(2)$ model, (IV.4). In contrast to dimensions three and higher, the Mermin–Wagner–Coleman theorem [36, 37] prohibits the presence of the massless Nambu–Goldstone particle with the linear dispersion in two dimensions. Thus, the $U(1)$ symmetry does not get broken even at low temperatures, while the system is in the superfluid phase, showing the algebraic long-range order. The effective Lagrangian of this superfluid phase is given by (II.12) when the vortex excitations are irrelevant.
@@
$\Upsilon_{\alpha}(L_{1}\times L_{2})=-T\frac{2}{\alpha^{2}}\frac{L_{2}}{L_{1}}\log\frac{Z_{\alpha}(L_{1}\times L_{2})}{Z_{0}(L_{1}\times L_{2})},$ (IV.11)
@@
$\Upsilon_{\alpha}(L_{1}\times L_{2})=\frac{T}{2\pi R_{\mathrm{ren}}^{2}}+O(\mathrm{e}^{-\#\frac{L_{1}}{L_{2}}})$ (IV.12)
@@
$\Upsilon_{\pi}(L_{1}\times L_{2})=\frac{T}{2\pi R_{\mathrm{ren}}^{2}}-\frac{2T}{\pi^{2}}\frac{L_{2}}{L_{1}}\ln 2+O(\mathrm{e}^{-\#\frac{L_{1}}{L_{2}}}).$ (IV.13)
@@
Equation (IV.12) shows that the helicity modulus $\Upsilon_{\alpha}$ dictates the renormalized compact boson radius directly for $0<\alpha\ll\pi$ when we can neglect the higher-winding contributions of $O(\mathrm{e}^{-\#L_{1}/L_{2}})$. This approximation becomes valid when $L_{1}/L_{2}\gg 1$. [^7] Ideally, it would be desired to take the limit $L_{1}/L_{2}\to\infty$ for completely eliminating higher-winding contributions. However, we should note that we compute $\Upsilon_{\alpha}$ out of $Z_{\alpha}/Z_{0}$, and this partition-function ratio itself becomes exponentially small as $L_{1}/L_{2}\to\infty$ according to (II.15). Thus, we should vary the ratio $L_{1}/L_{2}$ within the regime where $Z_{\alpha}/Z_{0}$ is computable with good accuracy. Moreover, as the higher-winding contributions are exponentially suppressed, we do not need to take extremely large $L_{1}/L_{2}$ in practice. This relationship becomes important when we apply the Nelson–Kosterlitz criterion [70] to obtain the BKT transition temperature $T_{\mathrm{BKT}}$, because it compares the helicity modulus $\Upsilon_{\alpha}$ with $2T/\pi$ to identify the marginality crossing point $R_{\mathrm{ren}}^{2}=1/4$. We note that this aspect-ratio dependence exists also for the helicity modulus defined by the local correlation function (or $\Upsilon_{\alpha}$ with the $\alpha\to 0$ limit), and see Refs. [71, 72], which study the aspect-ratio dependence of the helicity modulus. Although these higher-winding contributions turn out to be negligibly small within the numerical accuracy of this paper, [^8] For the $L\times L$ lattice, its contribution is already $O(10^{-4})$, and it becomes $O(10^{-9})$ for the $2L\times L$ lattice. we perform the computation not only at the $L\times L$ lattice but also at the $2L\times L$ lattice to suppress them.
@@
Figure 7 shows the helicity modulus $\Upsilon_{\alpha}(L\times L)$ evaluated at several twist angles, $\alpha=\pi,\pi/2,\pi/4,\pi/6$, for fixed $L=32$. We observe that the twist-angle dependence is quite tiny for $\Upsilon_{\alpha=\pi/2,\pi/4,\pi/6}$, which is consistent with the theoretical formula (IV.12) as its $\alpha$-dependence appears only in the subleading higher-winding contributions $O(\mathrm{e}^{-\#L_{1}/L_{2}})$. Furthermore, $\Upsilon_{\alpha=\pi}$ is significantly lower than the others on the low-temperature regions, which also matches with the theoretical expectation (IV.13). These helicity moduli at finite $\alpha$, obtained directly from the twisted partition functions, are found to be consistent with the recent Monte Carlo simulation results in Ref. [73].
@@
Figure 9: $T^{*}(L)$ as a function of $1/(\ln bL)^{2}$ for varying system sizes with the $2L\times L$ lattices: (a) $\alpha=\pi$, (b) $\alpha=\pi/6$. For $\alpha=\pi$, we determined $T^{*}(L)$ by using the modified Nelson–Kosterlitz criterion based on (IV.13). The results of different volumes follow the universal curve as a function of $1/(\ln bL)^{2}$, and we extract the BKT transition temperature by extrapolating $L\to\infty$.
Figure 9 shows $T^{*}(L)$ against $1/(\ln bL)^{2}$ at $\alpha=\pi$ and $\alpha=\pi/6$ evaluated on the $2L\times L$ lattices. When $\alpha=\pi$, $T^{*}(L)$ is determined based on Eq. (IV.13), not on the standard Nelson–Kosterlitz formula (IV.12), while the standard one is used for $\alpha=\pi/6$. In both cases, we simply neglect the correction from the higher-winding contributions as they are sufficiently small for the $2L\times L$ lattice with the current precision. The result is
$T_{\mathrm{BKT}}=\left\{\begin{array}[]{cl}0.8928(2)&(\text{from the $\alpha=\pi$ TRG}),\\ 0.8927(5)&(\text{from the $\alpha=\pi/6$ TRG}).\end{array}\right$. (IV.17)
@@ V Conclusions and Discussions
One of the potentially interesting applications is to consider the generalized $O(2)$ model, which extends the lattice action (IV.4) of the usual $O(2)$ model as [81, 82]
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available at Ref. [89].